OpenStudy (anonymous):

If tan a = 0.4750, find the value of a to the nearest minute

OpenStudy (therealmeeeee):

This is a example cot theta=0.7659 tan theta=1/0.7659 =1.3057 tan theta=1.3057 Use mathematical tables(which i don't have) to find the value of theta.

OpenStudy (wolf1728):

I think you have to determine the arc tangent of a to find a.

OpenStudy (anonymous):

So it would be tan-1(0.4750) = .4434483364

OpenStudy (wolf1728):

I think the answer is 25.408 Degrees then you have to convert the decimal portion to minutes of arc.

OpenStudy (anonymous):

As @wolf1728 said, you have to calculate:\[a=\arctan(0.7659)\] Then you'll need to convert the answers, which will either be in degrees or radians, to degrees and minutes.

OpenStudy (anonymous):

How do I do that

OpenStudy (wolf1728):

25.408 Degrees multiply .408 times 60

OpenStudy (anonymous):

So 25.408 is the nearest minute

OpenStudy (wolf1728):

No .408 x 60 = 24.48 minutes

OpenStudy (anonymous):

But how you got (.408) is it .4434483364

OpenStudy (wolf1728):

I didn't post .4434483364 Basically all I said was the arc tangent of .4750 is 25.408 Degrees

OpenStudy (anonymous):

Yea but I'm asking where you got .4750 from

OpenStudy (anonymous):

@wolf1728 got 25.408° which is the same as saying 25° +0.408°, now let's convert the 0.408° to minutes:\[\frac{ 0.408° }{ 1 }∙\frac{ 60 minutes }{ 1°}=answer\]Do you see now how to get minutes?

OpenStudy (anonymous):

See how the degrees cancel and leave you with an answer in terms of minutes?

OpenStudy (anonymous):

Oh now I get it